OpenStudy (acxbox22):

@gorechika12 please help ;)

OpenStudy (anonymous):

plese dont waste my time with a stupd question

OpenStudy (acxbox22):

calculus BC help

pooja195 (pooja195):

You must be famous acx question plz

OpenStudy (acxbox22):

can you please help me find the derivative of \[f(x)=(3x^2-4)^5\]

OpenStudy (acxbox22):

i know others can help me but you have proven yourself as a great helper

OpenStudy (anonymous):

whch one? all 3? soooo ez

OpenStudy (asad786):

who is that exactly

pooja195 (pooja195):

Do it than

pooja195 (pooja195):

then

OpenStudy (acxbox22):

its only one question

OpenStudy (misty1212):

chain chain chain... chain of rules

OpenStudy (anonymous):

u can strt with chain rule

OpenStudy (acxbox22):

can you lead me through it

OpenStudy (anonymous):

yes first you take out the 5 and reduce the powr to 4 d/dx of outside then inside

OpenStudy (misty1212):

the derivative of something to the power of five is five times something to the power of four, times the derivative of something

OpenStudy (acxbox22):

yes can you help me apply it to this problem?

OpenStudy (anonymous):

5(3x^2-4)^4 x 6x

OpenStudy (misty1212):

\[\frac{d}{dx}[\text{whatever}^5]=5\text{whatever}^4\times \frac{d}{dx}\text{whatever}\]

OpenStudy (acxbox22):

okay i get a little bit.. can you lead me to the answer please

OpenStudy (anonymous):

there is a program u can google wolfram u can teach urself. even tho i can do it 1000000000 timz bettr

OpenStudy (acxbox22):

the answer is \[y=x^22^x+\pi^2\] but can u explain it please

OpenStudy (anonymous):

r u stupid? where did pi come from?

OpenStudy (acxbox22):

the math book doesnt lie

OpenStudy (anonymous):

i correcd many math books at my age so yes they do

OpenStudy (acxbox22):

okay then can you please lead me to the correct answer i want to go to sleep

OpenStudy (acxbox22):

:( fine